OpenStudy (anonymous):

How to solve 39.50x + 31y= 578?

OpenStudy (anonymous):

solve for what???

OpenStudy (anonymous):

this is the question: Seventeen members of Troy’s family are going to an amusement park during the week of their family reunion. Adult tickets cost $39.50 each and child tickets cost $31 each. If the total cost for admission to the part is $578, how many adults went to the park? (3 points)

OpenStudy (anonymous):

so solving for x!

OpenStudy (anonymous):

39.50x=578-31y x= (578-31y)/39.50

OpenStudy (anonymous):

you get this??

OpenStudy (anonymous):

uhm can you help me solve that equation too , a little confused on where to start

OpenStudy (anonymous):

this is my answer choice, 6 8 9 11

OpenStudy (anonymous):

i m getting back you please i need some time dont mind it

OpenStudy (anonymous):

i got 6 adults and 11 children :) let me know